# Primal-Dual iLQR

### 05 Mar 2024 v2

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal-dual SQP problem. As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections, improving on existing direct multiple shooting approaches such as GNMS [3] and FDDP [6].
@@ II-C Contributions
This paper presents a self-contained description of the specialization of NPSQP to unconstrained discrete-time optimal control problems. As compared to GNMS and FDDP, our method has the advantages of being known to be globally convergent and to not impede superlinear local convergence. Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners.
This framework can be extended to also handle constrained problems, using augmented Lagrangian [5], interior point [12], or SQP [13] methods. ALTRO would benefit in using the approach described here to avoid the control dimension expansion it currently performs (resulting in performance losses). Moreover, acados and FATROP would benefit in using the globalization strategy described here, which does not require second order corrections.
@@ IV Algorithm Derivation
V Algorithm Overview
In this section, we provide some pseudocode for the proposed algorithm. For the sake of simplicity, we exclude regularization from the code here; we also exclude tolerance considerations when performing zero-checks.
@@ V Algorithm Overview
VI Benchmarks
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. This seems to be the problem used in [16]. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step. A JAX implementation of our algorithm takes 46ms to finish solving this problem. Below, you can find a visual description of the problem and solution, as well as solver logs.
@@ VI Benchmarks
VII Conclusion
This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm. This method improves on earlier direct multiple shooting methods (such as GNMS and FDDP), specifically by guaranteeing global convergence and not impeding local superlinear convergence, under the assumptions described in [4].
Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton. Moreover, it allows that the evaluation of the dynamics and derivatives be computed in parallel.
Our algorithm can also be used to significantly speed up ALTRO, one of the state of the art open source software packages for solving constrained discrete-time optimal control problems, by removing the control dimension expansion occurring therein. Our method also removes the need to handle second order corrections, which would simplify and potentially speed up FATROP.

### 13 Mar 2024 v3

@@ I Abstract
We introduce a new algorithm for solving unconstrained discrete-time optimal control problems. Our method follows a direct multiple shooting approach, and consists of applying the SQP method together with an $\ell_{2}$ augmented Lagrangian primal-dual merit function. We use the LQR algorithm to efficiently solve the primal component of the Newton-KKT system, and use a dual LQR backward pass to solve its dual component. We also present a new parallel algorithm for solving the dual component of the Newton-KKT system in $O(\log(N))$ parallel time, where $N$ is the number of stages. Combining it with [17], we are able to solve the full Newton-KKT system in $O(\log(N))$ parallel time. The remaining parts of our method have constant parallel time complexity per iteration. Therefore, this paper provides, for the first time, a practical, highly parallelizable (for example, with a GPU) method for solving nonlinear discrete-time optimal control problems. As our algorithm is a specialization of NPSQP [4], it inherits its generic properties, including global convergence, fast local convergence, and the lack of need for second order corrections or dimension expansions, improving on existing direct multiple shooting approaches such as acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6].
@@ II-C Contributions
This paper introduces the Primal-Dual iLQR algorithm, which consists of a specialization of the generic NPSQP algorithm to the case of unconstrained discrete-time optimal control problems.
We also introduce a new method for parallelizing the computation of the dual solution of the Newton-KKT systems, with parallel time complexity $O(\log(N)\log(n))$, by framing the dual LQR backward pass as an associative scan [18]. This can be combined with the method of [17] to solve the primal-dual Newton-KKT system in parallel time complexity logarithmic in $N$.
Our algorithm improves on earlier direct multiple shooting methods, specifically by guaranteeing global convergence and not impeding local superlinear convergence (even without second order corrections), under the assumptions described in [4].
@@ IV Algorithm Derivation
V Parallel Dual LQR Backward Pass with Associative Scans
Associative scans are a common parallelization mechanism used in functional programming, first introduced in [18]. They were used in [17] to derive a simple method for solving (primal) LQR problems in $O(\log(m)+\log(N)\log(n))$ parallel time, where $N,n,m$ are respectively the number of stages, states, and controls.
In this section, we give a brief description of what associative scans are, and how they can also be used to derive a similar method that is equivalent to the dual LQR backward pass, while only taking $O(\log(N)\log(n))$ parallel time. Note that the $\log(n)$ factor here comes from the parallel time complexity of taking the product of two matrices of dimension $n$.
Given a set $\mathcal{X}$, a function $f:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{X}$ is said to be associative if $\forall a,b,c\in\mathcal{X},f(f(a,b),c)=f(a,f(b,c))$. The forward associative scan operation $S_{f}(x_{1},\ldots,x_{n};f)$ can be recursively defined by $S_{f}(x_{1};f)=x_{1}$ and $S_{f}(x_{1},\ldots,x_{i+1};f)=(y_{1},\ldots,y_{i},f(y_{i},x_{i+1}))$, where $y_{1},\ldots,y_{i}=S_{f}(x_{1},\ldots,x_{i};f)$. Similarly, the reverse associative scan operation $S_{r}(x_{1},\ldots,x_{n};f)$ can be recursively defined by $S_{r}(x_{1};f)=x_{1}$ and $S_{r}(x_{1},\ldots,x_{i+1};f)=(f(x_{1},y_{2}),y_{2},\ldots,y_{i+1})$, where $y_{2},\ldots,y_{i+1}=S_{r}(x_{2},\ldots,x_{i+1};f)$. [18] provides a method for performing associative scans of $N$ elements in parallel time $O(\log(N))$.
As shown in eq. 47, the dual solutions $\Delta\lambda_{i}$ of the Newton-KKT system eq. 10 can be obtained by applying a backward pass starting at $\Delta\lambda_{N}=Q_{N}\Delta x_{N}+l_{N}$, with recursion rule $\Delta\lambda_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+A_{i}^{T}\Delta\lambda_{i+1}+l_{i}^{x}$. We can simplify the notation by defining, for $i\in\{0,\ldots,N-1\}$, $M_{i}=A_{i}^{T}$ and $m_{i}=Q_{i}^{xx}\Delta x_{i}+Q_{i}^{xu}\Delta u_{i}+l_{i}^{x}$, so that the recursion rule becomes simply $\Delta\lambda_{i}=M_{i}\Delta\lambda_{i+1}+m_{i}$.
Moreover, we define $m_{N}=Q_{N}\Delta x_{N}+l_{N}$ and $M_{N}=I$.
Letting $\mathcal{X}=\mathbb{R}^{n}\times\mathbb{R}^{n\times n}$ and letting $f:\mathcal{X}\rightarrow\mathcal{X}$ be defined by $f((a,B),(c,D))=(a+Bc,BD)$, we claim that $f$ is associative and that $\Delta\lambda_{0},\ldots,\Delta\lambda_{N}$ can be obtained by doing a reverse associative scan with respect to $f$ on $(m_{0},M_{0}),\ldots,(m_{N-1},M_{N-1}),(m_{N},M_{N})$.
The associative property is simple to verify:
$f(f((a,B),(c,D)),(e,F))$ (48)
$=$ $f((a+Bc,BD),(e,F))$
$=$ $(a+Bc+BDe,BDF)$
$=$ $(a+B(c+De),BDF)$
$=$ $f((a,B),(c+De,DF))$
$=$ $f((a,B),f((c,D),(e,F))).$
Let $F$ denote the operator that extracts the first element from its input sequence. We can show by induction that
$(\Delta\lambda_{i},M_{i}\cdots M_{N})=F(S_{r}((m_{i},M_{i})),\ldots,(m_{N},M_{N});f).$ (49)
This is clearly true for $i=N$, as
$S_{r}((m_{N},M_{N});f)=(m_{N},M_{N})=(\Delta\lambda_{N},M_{N}).$ (50)
For $i<N$, by definition of $S_{r}$ and the induction hypothesis,
$F(S_{r}((m_{i},M_{i}),\ldots,(m_{N},M_{N});f))$ (51)
$=$ $f((m_{i},M_{i}),F(S_{r}((m_{i+1},M_{i+1}),\ldots,(m_{N},M_{N});f)))$
$=$ $f((m_{i},M_{i}),(\Delta\lambda_{i+1},M_{i+1}\cdots M_{N}))$
$=$ $(m_{i}+M_{i}\Delta\lambda_{i+1},M_{i}\cdots M_{N})$
$=$ $(\Delta\lambda_{i},M_{i}\cdots M_{N}).$
VI Algorithm Overview
In this section, we provide some pseudocode for the sequential version of the proposed algorithm. For the sake of simplicity, we exclude regularization from the code here; we also exclude tolerance considerations when performing zero-checks.
@@ VI Algorithm Overview
VII Benchmarks
Below, we include an example run of our algorithm on on a quad-pendulum problem, which we got from the Trajax repository on GitHub and modified to treat constraints as high-penalty costs (using a gain of $100$). We also increased the final state cost gains on the positions to $1000$. This seems to be the problem used in [16]. Note that we solve a single primal-dual LQR problem for each iteration, and have to evaluate the user model once per line search step. A JAX implementation of our algorithm takes 46ms to finish solving this problem (on the CPU of an M2 MacBook Air). Below, you can find a visual description of the problem and solution, as well as solver logs.
@@ VII Benchmarks
VIII Conclusion
This paper introduces the Primal-Dual iLQR algorithm, a specialization of the generic NPSQP algorithm.
Due to relying mostly on generic optimization techniques and not needing a tailored theoretical analysis, the proposed method has a simple mathematical derivation, making it easily accessible to practicioners. Moreover, it inherits all the desirable properties that NPSQP has been shown to have, such as not impeding local superlinear convergence, global convergence, and the lack of need for second order corrections. None of acados [13], ALTRO [5], GNMS [3], FATROP [12], and FDDP [6] combine all of these properties.
The proposed algorithm has sequential time complexity $O(N(n+m)^{3})$, where $N,n,m$ are respectively the number of stages, states, and controls of the given problem, and parallel time complexity $O(\log(m)+\log(N)\log(n)$. Note that a naïf application of the NPSQP algorithm would have sequential time complexity $O(N^{3}(n+m)^{3})$ and parallel time complexity $O(\log(N)+\log(n+m))$ (although requiring a much larger number of workers), as the Newton-KKT systems have dimension $(N+1)(n+m)$. This is due to the fact that inverting a matrix of dimension $d$ takes $O(d^{3})$ sequential time and $O(\log(d))$ parallel time.
This framework can easily be extended to also handle constrained problems, ideally using an augmented Lagrangian [5] or interior point [12] method; using SQP [13] to handle general nonlinear constraints would yield Newton-KKT systems that are fundamentally different from the ones we consider, and would therefore require substantial changes to our method.
Our method is substantially easier to warm start as compared to single-shooting methods such as DDP, iLQR, or Stagewise Newton. Moreover, as it never performs dynamics rollouts, it allows that the evaluation of the dynamics and derivatives be computed in parallel, resulting in $O(1)$ parallel time per line search iteration.